OpenStudy (anonymous):

If f is continuous and interval[0,2] f(x) dx =6, find interval [0,pi/2] 3f(2sinx)cosx dx

OpenStudy (anonymous):

I assume by that you mean integral, not interval, right?

OpenStudy (anonymous):

well i mean the upper limit is 2 and lower is 0 like \[\int\limits_{0}^{2}f (x)dx\] , find \[\int\limits_{0}^{\pi/2} 3f(2sinx)cosxdx\]

OpenStudy (anonymous):

So, yes, you mean integral.

OpenStudy (anonymous):

Yeah, that's what I thought. Let \[x = 2sin(\theta) \rightarrow dx = 2cos(\theta)d\theta\] then we have that \[\theta = \sin^{-1}(\frac{x}{2})\] so at the endpoints, \[\theta = \sin^{-1}(0) = 0, \theta = \sin^{-1}(1) = \frac{\pi}{2} \] so the integral becomes \[\int_0^{\frac{\pi}{2}} f(2\sin(\theta))\cdot 2\cos(\theta) d\theta\] The substitution doesn't change the value of the integral, so we have that \[2\int_0^{\frac{\pi}{2}} f(2\sin(\theta))\cdot\cos(\theta) d\theta = 6\] so \[\int_0^{\frac{\pi}{2}} f(2\sin(\theta))\cdot\cos(\theta) d\theta = 3 \] The question asks for \[3\int_0^{\frac{\pi}{2}} f(2\sin(\theta))\cdot\cos(\theta) d\theta \] which is apparently 3*3 = 9.

OpenStudy (anonymous):